# CAST: Component-Aligned 3D Scene Reconstruction from an RGB Image

### v1

@@ 2.3 Physics-Aware 3D Modeling
Generating physically plausible 3D assets is crucial for ensuring realism and functionality in applications such as animation, gaming, and robotics. While recent 3D generative models excel at creating visually realistic objects, they often fall short in achieving physical plausibility. To address this limitation, physics-aware 3D generative models have been developed to integrate physical principles into the generation process. Some methods use soft-body simulation to animate 3D Gaussians [81, 94], or generate articulated objects with physics-based penalties [49], while others ensure self-supporting structures through rigid-body simulation [55, 54, 13] or FEM [29, 82]. These methods leverage offline or online physical simulations to check the physical validity of generated shapes and in turn guide generation. However, these approaches are typically confined to individual objects, overlooking the mutual influences between multiple objects within a scene.
@@
Incorporating physical constraints into scene synthesis is much more challenging due to the inclusion of more complex relationships, e.g., inter-object contact. [85] integrates constraints like object collisions, room layout, and object reachability into their scene-level generation pipeline. However, it is limited to indoor scene synthesis and relies on a closed-vocabulary database to perform shape retrieval. [58] addresses the issue of physical implausibility in multi-view neural reconstruction. It leverages both differentiable rendering and physical simulation to learn implicit representations. However, it requires multi-view images as input, focuses on individual objects, and primarily addresses only stability (simulating the dropping of objects). In contrast, our method operates in an open-vocabulary setting and requires only a single input image. Furthermore, it accounts for more complex inter-object relationships, particularly support and contact, making it more versatile and applicable across diverse scenarios.
3 Overview
@@
Scene-level reconstruction from a single image is a fundamental challenge in computer graphics, with broad applications in animation, virtual reality, and interactive gaming. Unlike object-level reconstruction, which focuses on isolated objects, scene-level reconstruction emphasizes the arrangement and relations of multiple entities under realistic (or stylized) physics. By capturing per-object structures, spatial relationships, and contextual cues, this holistic approach enables more immersive experiences, compelling narratives, and efficient workflows—benefits that surpass those of single-object reconstructions. Although previous methodologies have explored feed-forward pipelines or retrieval-based approaches using fixed 3D templates [45, 18], these methods often struggle to capture nuanced scene semantics and complex object relationships. To address these limitations, we propose a generation-driven scene “reincarnation” approach with emphasized object relations to construct high-fidelity, contextually consistent 3D environments from a single, unannotated RGB image whether sourced from real-world photography or synthetic data (see Fig. 2).
@@
A key insight of our method is the thorough object relation analysis of scene contextual information. First, we perform object segmentation to identify and localize constituent objects within the image. We then obtain preliminary geometric information, i.e., point clouds, and explore semantic and spatial relationships among objects. This contextual backbone informs our subsequent object-wise generation pipeline, ensuring that each reconstructed object retains not only its geometric fidelity but also its correct placement in the broader scene. Finally, we synthesize a coherent 3D environment that respects physical plausibility—achieving structurally sound layouts and realistic interactions among scene elements.
@@
Our research focuses on two primary objectives: to explore how generative models can effectively capture complex inter-object relationships in order to produce realistic, scene-level reconstructions from a single image; and to identify strategies for integrating geometric cues and contextual information that maximize accuracy and plausibility in 3D reconstructions. Through this investigation, we demonstrate that generative methods provide a more flexible and robust alternative to traditional feed-forward and retrieval-based techniques. These methods allow for fine-grained control over both object-level details and global scene composition, thus streamlining content creation pipelines for animation, game development, and other fields requiring accurate, visually compelling 3D models. This work highlights the advantages of a generation-centric framework and lays the groundwork for future advancements in scene-level 3D reconstruction. It also underscores the growing importance of context-driven approaches in bridging the gap between 2D imagery and immersive, interactive virtual environments.
@@
To facilitate comprehensive scene reconstruction from a single image, we first perform an extensive semantic extraction that provides a robust foundation for subsequent processing. Specifically, we employ Florence-2 [80] to identify objects, generate their descriptions, and localize each object with bounding boxes. We then leverage GPT-4v [1] to filter out spurious detections and isolate meaningful constituent objects, allowing for open-vocabulary object identification that is not constrained by predefined categories. Next, we use GroundedSAM-v2 [64] to produce a refined segmentation mask $\{\bm{M}_{i}\}$ for each labeled object $\{\bm{o}_{i}\}$, thereby obtaining both precise object boundaries and corresponding occlusion masks, which play a crucial auxiliary role in the object generation stage. Apart from semantic cues, we also integrate geometric information by extracting a scene-level point cloud. Using MoGe [74], we generate pixel-aligned point clouds $\{\bm{q}_{i}\}$ for each object $\{\bm{o}_{i}\},i\in\{1,\dots,N\}$ and a global camera parameter in the scene coordinate system. This additional geometric data is subsequently matched to each object’s segmentation mask, providing a reliable structural reference for the final 3D scene reconstruction.
@@
Figure 3: Network design of our alignment generation model (Sec. 4.2), occlusion-aware object generation model (Sec. 4.1), and an illustrative figure of the texture generation model.
4 Perceptive 3D Instance Generation
@@
In the endeavor to reconstruct high-fidelity 3D scenes from single RGB images, a brute-force approach involves generating the entire scene mesh directly using techniques such as single-image depth estimation or diffusion priors. However, this method inherently struggles to manage occlusions, render invisible components, and accurately represent object relationships due to the complex and intertwined nature of real-world scenes. Instead of generating an entire scene mesh directly, our approach focuses on individual object generation and then arranges the objects via precise relational alignment, as illustrated in 3. This strategy offers several advantages: 1. focusing on individual objects ensures higher geometric fidelity and allows for detailed modeling, resulting in more accurate and visually appealing scene components. 2. operating within a canonical space ensures that generated assets adhere to standardized orientations and scales, seamlessly integrating with artist-defined coordinate systems and promoting consistency across digital content creation tools. 3. the modular approach supports various applications such as editing, rendering, and simulation, enabling independent manipulation of objects for greater flexibility and efficiency. By decomposing scene reconstruction into object-wise generation and alignment, our method improves asset quality and manageability while enhancing the overall coherence and functionality of the 3D environment. This approach addresses challenges like geometric precision and efficient post-processing, advancing single-image 3D scene generation.
@@
A critical aspect of our object generation process is the utilization of a large generative model to generate holistic and high-fidelity object meshes from partial image and point cloud observations. To do so, we first follow state-of-art native 3D generative models [79, 90, 92] to pre-train a large-scale 3D generative model conditioning on textual and image inputs. We build upon existing generative frameworks featuring 3DShape2VecSet representation [90, 92], which prioritize geometry generation by utilizing a Geometry Variational Autoencoder (VAE). This VAE framework encodes uniformly sampled surface point clouds into unordered latent codes and decodes these latent representations into Signed Distance Fields (SDFs). Formally, the VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are defined as:
$\bm{Z}=\mathcal{E}(\bm{X}),\ \mathcal{D}(\bm{Z},\bm{p})=\text{SDF}(\bm{p}),$ (1)
where $\bm{X}$ represents the sampled surface point cloud of the geometry, $\bm{Z}$ is the corresponding latent code, and $\text{SDF}(\bm{p})$ denotes the operation of querying the SDF value at point $\bm{p}$ for subsequent mesh extraction via marching cubes. To effectively incorporate image information into the geometry generation process, we employ DINOv2 [59] as our image encoder, following methodologies outlined in [90, 92, 79], The geometry latent diffusion model (LDM) is then formulated as:
$\epsilon_{\text{obj}}(\bm{Z}_{t};t,\bm{c})\rightarrow\bm{Z},$ (2)
where $\epsilon$ represents the diffusion transformer model, $\bm{Z}_{t}$ is noisy geometry latent code at timestep $t$, and $\bm{c}$ denotes the encoded image features from DINOv2. We follow the pre-training process of prior works [90, 92] and pre-train the base model on Objaverse [20]. Upon training, our generation model $\epsilon$ is capable of generating detailed 3D geometry solely based on image features.
4.1 Occlusion-aware 3D Object Generation
@@
$\bm{c}_{m}=\mathcal{E}_{\text{DINOv2}}(\bm{I}\odot\bm{M}),$ (3)
@@
Though our object generation model produces visually plausible meshes from input object images, it is challenging to generate pixel-aligned geometry due to the high-level nature of the encoded image condition $\bm{c}$ and the absence of pixel-wise supervision. We address this issue by additionally conditioning our object generation model on observed partial point clouds in canonical coordinates. This dual conditioning ensures that the generated geometries not only align visually with the input images but also accurately reflect their underlying scale, shape, and depth. During the conditioning training, we simulate real-world partial scans or estimated depth maps by rendering each 3D asset from multiple viewpoints, thereby obtaining corresponding RGB images, camera parameters, and ground-truth depth maps. These RGB images are then processed using advanced depth estimation techniques, including MoGe [74] and Metric3D [84], to produce an estimated depth map and then projected as partial point clouds. These point clouds are coarsely aligned with the original objects within the canonical $[-1,1]^{3}$ space.
@@
$\epsilon(\bm{Z}_{t};t,\bm{c},\bm{p}_{\text{disturb}})\rightarrow\bm{Z},$ (4)
@@
where the conditioning adaptation scheme is based on attention mechanism similar to [90, 92]. Additionally, to mimic real-world occlusions and missing data, we randomly mask sets of basic primitives—such as circles and rectangles—in the depth maps from various camera views. This results in partial point clouds with occluded and incomplete regions, further enhancing the model’s ability to handle imperfect inputs. A critical design choice in our approach is to maintain the alignment of partial point clouds with the geometry in our training data set. Unlike methods that apply random scaling, translation, or rotation to augmented point clouds, our aligned partial point clouds ensure that the generative model can more effectively conform to the input point clouds’s inherent structure. This alignment restricts the model to adhere closely to the actual shapes and scales of objects, thereby facilitating more precise and coherent 3D reconstructions. By conditioning on these well-aligned partial point clouds, our model achieves superior alignment both in overall size and local geometric details, resulting in high-quality and reliable 3D geometry generation.
4.2 Generative Alignment
@@
Each generated 3D object is within a normalized volume and assumes a canonical pose that may not be aligned with the image and scene space point cloud. This is because the image conditions use high-level features, such as DINOv2, to achieve better generalization. Ensuring that each object is correctly transformed and scaled to align with its presentation in the scene is crucial for scene composition. Though traditional alignment methods, such as Iterative Closest Point (ICP) [2, 6], can be employed, they often fail to account for semantic context, leading to frequent misalignments and diminished accuracy (see Fig. 9). Instead, we introduce an alignment generative model conditioned on the scene-space partial point cloud $\bm{q}\in\mathbb{R}^{N\times 3}$ and the canonical-space geometry latent code $\bm{Z}$. Formally, we define our alignment generator “AlignGen” as:
$\epsilon_{\text{align}}(\bm{p}_{t};t,\bm{q},\bm{Z})\rightarrow\bm{p},$ (5)
@@
where $\epsilon_{\text{align}}$ is a point cloud diffusion transformer, $\bm{p}\in\mathbb{R}^{N\times 3}$ is the transformed version of the scene-space partial point cloud to the canonical space, aligning with the generated object mesh. $\bm{Z}$ is the generated geometry latent of object corresponding to $\bm{p}$ from the object generation model. $\bm{p}_{t}$ is the noised version of $\bm{p}$ at timestep $t$. In essence, the generation model maps the scene-space partial point cloud $\bm{q}$ to $\bm{p}$ in the canonical $[-1,1]^{3}$ space, aligning it with the generated object mesh. We can subsequently recover the similarity transformation (i.e., scaling, rotation, and translation) from $\bm{q}$ and $\bm{p}$ using the Umeyama algorithm [72] as they are point-wise corresponded. This final step is numerically more stable than directly predicting transformation parameters.
@@
In practice, we employ distinct conditioning strategies for the input point cloud $\bm{q}$ and the geometric latent $\bm{Z}$. For $\bm{q}$, we concatenate the input point cloud with the diffusion sample $\bm{p}_{t}$ along the feature channel dimension, enabling the transformer architecture to learn explicit correspondences between the noisy canonical-frame partial cloud and the world-space partial cloud. For the geometric latent $\bm{Z}$, we apply a cross-attention mechanism to inject it into the point diffusion transformer. This approach ensures that the model effectively incorporates spatial and geometric relationships. Additionally, due to symmetry and replicated geometrical shapes, multiple valid $\bm{p}$ may exist for a given $\bm{q}$ and $\bm{Z}$. Our diffusion model addresses this by sampling multiple noise realizations and aggregating the resulting transformations, to select the most confident and coherent representations.
4.3 Iterative Generation Procedure

### v2

@@ 2.3. Physics-Aware 3D Modeling
Generating physically plausible 3D assets is crucial for ensuring realism and functionality in applications such as animation, gaming, and robotics. While recent 3D generative models excel at creating visually realistic objects, they often fall short in achieving physical plausibility. To address this limitation, physics-aware 3D generative models have been developed to integrate physical principles into the generation process. Some methods use soft-body simulation to animate 3D Gaussians (Xie et al., 2024; Zhong et al., 2025), or generate articulated objects with physics-based penalties (Liu et al., 2023b), while others ensure self-supporting structures through rigid-body simulation (Mezghanni et al., 2021, 2022; Chen et al., 2024b) or FEM (Guo et al., 2024; Xu et al., 2024). These methods leverage offline or online physical simulations to check the physical validity of generated shapes and in turn guide generation. However, these approaches are typically confined to individual objects, overlooking the mutual influences between multiple objects within a scene.
@@
Incorporating physical constraints into scene synthesis is much more challenging due to the inclusion of more complex relations, e.g., inter-object contact. Yang et al. (2024a) integrates constraints like object collisions, room layout, and object reachability into their scene-level generation pipeline. However, it is limited to indoor scene synthesis and relies on a closed-vocabulary database to perform shape retrieval. Ni et al. (2024) addresses the issue of physical implausibility in multi-view neural reconstruction. It leverages both differentiable rendering and physical simulation to learn implicit representations. However, it requires multi-view images as input, focuses on individual objects, and primarily addresses only stability (simulating the dropping of objects). In contrast, our method operates in an open-vocabulary setting and requires only a single input image. Furthermore, it accounts for more complex inter-object relations, particularly support and contact, making it more versatile and applicable across diverse scenarios.
3. Overview
@@
Scene-level reconstruction from a single image is a fundamental challenge in computer graphics, with broad applications in animation, virtual reality, and interactive gaming. Unlike object-level reconstruction, which focuses on isolated objects, scene-level reconstruction emphasizes the arrangement and relations of multiple entities under realistic (or stylized) physics. By capturing per-object structures, spatial relations, and contextual cues, this holistic approach enables more immersive experiences, compelling narratives, and efficient workflows—benefits that surpass those of single-object reconstructions. Although previous methodologies have explored feed-forward pipelines or retrieval-based approaches using fixed 3D templates (Liu et al., 2022; Dai et al., 2024), these methods often struggle to capture nuanced scene semantics and complex object relations. To address these limitations, we propose a generation-driven scene reconstruction approach with emphasized object relations to construct high-fidelity, contextually consistent 3D environments from a single, unannotated RGB image whether sourced from real-world photography or synthetic data (see Fig. 2).
@@
A key insight of our method is the thorough object relation analysis of scene contextual information. First, we perform object segmentation to identify and localize constituent objects within the image. We then obtain preliminary geometric information, i.e., point clouds, and explore semantic and spatial relations among objects. This contextual backbone informs our subsequent object-wise generation pipeline, ensuring that each reconstructed object retains not only its geometric fidelity but also its correct placement in the broader scene. Finally, we synthesize a coherent 3D environment that respects physical plausibility—achieving structurally sound layouts and realistic interactions among scene elements.
@@
Our research focuses on two primary objectives: to explore how generative models can effectively capture complex inter-object relations in order to produce realistic, scene-level reconstructions from a single image; and to identify strategies for integrating geometric cues and contextual information that maximize accuracy and plausibility in 3D reconstructions. Through this investigation, we demonstrate that generative methods provide a more flexible and robust alternative to traditional feed-forward and retrieval-based techniques. These methods allow for fine-grained control over both object-level details and global scene composition, thus streamlining content creation pipelines for animation, game development, and other fields requiring accurate, visually compelling 3D models. This work highlights the advantages of a generation-centric framework and lays the groundwork for future advancements in scene-level 3D reconstruction. It also underscores the growing importance of context-driven approaches in bridging the gap between 2D imagery and immersive, interactive virtual environments.
@@
To facilitate comprehensive scene reconstruction from a single image, we first perform an extensive semantic extraction that provides a robust foundation for subsequent processing. Specifically, we employ Florence-2 (Xiao et al., 2024) to identify objects, generate their descriptions, and localize each object with bounding boxes. We then leverage GPT-4v (Achiam et al., 2023) to filter out spurious detections and isolate meaningful constituent objects, allowing for open-vocabulary object identification that is not constrained by predefined categories. Next, we use GroundedSAM-v2 (Ren et al., 2024) to produce a refined segmentation mask $\{\bm{M}_{i}\}$ for each labeled object $\{\bm{o}_{i}\}$, thereby obtaining both precise object boundaries and corresponding occlusion masks, which play a crucial auxiliary role in the object generation stage. Apart from semantic cues, we also integrate geometric information by extracting a scene-level point cloud. Using MoGe (Wang et al., 2024b), we generate pixel-aligned point clouds $\{\bm{q}_{i}\}$ for each object $\{\bm{o}_{i}\},i\in\{1,\dots,N\}$ and a global camera parameter in the scene coordinate system. This additional geometric data is subsequently matched to each object’s segmentation mask, providing a reliable structural reference for the final 3D scene reconstruction.
@@
Figure 3. Network design of our alignment generation model (Sec. 4.2), occlusion-aware object generation model (Sec. 4.1), and an illustrative figure of the texture generation model.
4. Perceptive 3D Instance Generation
@@
In the endeavor to reconstruct high-fidelity 3D scenes from single RGB images, a brute-force approach involves generating the entire scene mesh directly using techniques such as single-image depth estimation or diffusion priors. However, this method inherently struggles to manage occlusions, render invisible components, and accurately represent object relations due to the complex and intertwined nature of real-world scenes. Instead of generating an entire scene mesh directly, our approach focuses on individual object generation and then arranges the objects via precise relational alignment, as illustrated in 3. This strategy offers several advantages: 1. focusing on individual objects ensures higher geometric fidelity and allows for detailed modeling, resulting in more accurate and visually appealing scene components. 2. operating within a canonical space ensures that generated assets adhere to standardized orientations and scales, seamlessly integrating with artist-defined coordinate systems and promoting consistency across digital content creation tools. 3. the modular approach supports various applications such as editing, rendering, and simulation, enabling independent manipulation of objects for greater flexibility and efficiency. By decomposing scene reconstruction into object-wise generation and alignment, our method improves asset quality and manageability while enhancing the overall coherence and functionality of the 3D environment. This approach addresses challenges like geometric precision and efficient post-processing, advancing single-image 3D scene generation.
@@
A critical aspect of our object generation process is the utilization of a large generative model to generate holistic and high-fidelity object meshes from partial image and point cloud observations. To do so, we first follow state-of-art native 3D generative models (Xiang et al., 2024; Zhang et al., 2023, 2024a) to pre-train a large-scale 3D generative model conditioning on textual and image inputs.
We build upon existing generative frameworks featuring 3DShape2VecSet representation (Zhang et al., 2023, 2024a), which prioritize geometry generation by utilizing a Geometry Variational Autoencoder (VAE). This VAE framework encodes uniformly sampled surface point clouds into unordered latent codes and decodes these latent representations into Signed Distance Fields (SDFs). Formally, the VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are defined as:
(1) $\bm{Z}=\mathcal{E}(\bm{X}),\ \mathcal{D}(\bm{Z},\bm{p})=\text{SDF}(\bm{p}),$
where $\bm{X}$ represents the sampled surface point cloud of the geometry, $\bm{Z}$ is the corresponding latent code, and $\text{SDF}(\bm{p})$ denotes the operation of querying the SDF value at point $\bm{p}$ for subsequent mesh extraction via marching cubes. To effectively incorporate image information into the geometry generation process, we employ DINOv2 (Oquab et al., 2023) as our image encoder, following methodologies outlined in Zhang et al. (2023, 2024a); Xiang et al. (2024), The geometry latent diffusion model (LDM) is then formulated as:
(2) $\epsilon_{\text{obj}}(\bm{Z}_{t};t,\bm{c})\rightarrow\bm{Z},$
where $\epsilon$ represents the diffusion transformer model, $\bm{Z}_{t}$ is noisy geometry latent code at timestep $t$, and $\bm{c}$ denotes the encoded image features from DINOv2. We follow the pre-training process of prior works (Zhang et al., 2023, 2024a) and pre-train the base model on Objaverse (Deitke et al., 2023). Upon training, our generation model $\epsilon$ is capable of generating detailed 3D geometry solely based on image features.
4.1. Occlusion-aware 3D Object Generation
@@
(3) $\bm{c}_{m}=\mathcal{E}_{\text{DINOv2}}(\bm{I}\odot\bm{M}),$
@@
Though our object generation model produces visually plausible meshes from input object images, it is challenging to generate pixel-aligned geometry due to the high-level nature of the encoded image condition $\bm{c}$ and the absence of pixel-wise supervision. We address this issue by additionally conditioning our object generation model on observed partial point clouds in canonical coordinates. This dual conditioning ensures that the generated geometries not only align visually with the input images but also accurately reflect their underlying scale, shape, and depth. During the conditioning training, we simulate real-world partial scans or estimated depth maps by rendering each 3D asset from multiple viewpoints, thereby obtaining corresponding RGB images, camera parameters, and ground-truth depth maps. These RGB images are then processed using advanced depth estimation techniques, including MoGe (Wang et al., 2024b) and Metric3D (Yang et al., 2024b), to produce an estimated depth map and then projected as partial point clouds. To ensure scale consistency, we align estimated depth maps from MoGe and Metric3D with ground-truth depth maps by scaling and shifting them based on the median and median absolute deviation of valid depth values. The resulting point clouds are then normalized to a canonical [-1, 1]3 space to ensure consistent spatial representation for coarse object alignment.
@@
(4) $\epsilon(\bm{Z}_{t};t,\bm{c},\bm{p}_{\text{disturb}})\rightarrow\bm{Z},$
@@
where the conditioning adaptation scheme is based on attention mechanism similar to Zhang et al. (2023, 2024a). Additionally, to mimic real-world occlusions and missing data, we randomly mask sets of basic primitives—such as circles and rectangles—in the depth maps from various camera views. This results in partial point clouds with occluded and incomplete regions, further enhancing the model’s ability to handle imperfect inputs. A critical design choice in our approach is to maintain the alignment of partial point clouds with the geometry in our training data set. Unlike methods that apply random scaling, translation, or rotation to augmented point clouds, our aligned partial point clouds ensure that the generative model can more effectively conform to the input point clouds’s inherent structure. This alignment restricts the model to adhere closely to the actual shapes and scales of objects, thereby facilitating more precise and coherent 3D reconstructions. By conditioning on these well-aligned partial point clouds, our model achieves superior alignment both in overall size and local geometric details, resulting in high-quality and reliable 3D geometry generation.
4.2. Generative Alignment
@@
Each generated 3D object is within a normalized volume and assumes a canonical pose that may not be aligned with the image and scene space point cloud. This is because the image conditions use high-level features, such as DINOv2, to achieve better generalization. Ensuring that each object is correctly transformed and scaled to align with its presentation in the scene is crucial for scene composition. Though traditional alignment methods, such as Iterative Closest Point (ICP) (Arun et al., 1987; Best, 1992), can be employed, they often fail to account for semantic context, leading to frequent misalignments and diminished accuracy (see Fig. 9). Instead, we introduce an alignment generative model conditioned on the scene-space partial point cloud $\bm{q}\in\mathbb{R}^{N\times 3}$ and the canonical-space geometry latent code $\bm{Z}$. Formally, we define our alignment generator “AlignGen” as:
(5) $\epsilon_{\text{align}}(\bm{p}_{t};t,\bm{q},\bm{Z})\rightarrow\bm{p},$
@@
where $\epsilon_{\text{align}}$ is a point cloud diffusion transformer, $\bm{p}\in\mathbb{R}^{N\times 3}$ is the transformed version of the scene-space partial point cloud to the canonical space, aligning with the generated object mesh. $\bm{Z}$ is the generated geometry latent of object corresponding to $\bm{p}$ from the object generation model. $\bm{p}_{t}$ is the noised version of $\bm{p}$ at timestep $t$. In essence, the generation model maps the scene-space partial point cloud $\bm{q}$ to $\bm{p}$ in the canonical $[-1,1]^{3}$ space, aligning it with the generated object mesh. We can subsequently recover the similarity transformation (i.e., scaling, rotation, and translation) from $\bm{q}$ and $\bm{p}$ using the Umeyama algorithm (Umeyama, 1991) as they are point-wise corresponded. This final step is numerically more stable than directly predicting transformation parameters.
@@
In practice, we employ distinct conditioning strategies for the input point cloud $\bm{q}$ and the geometric latent $\bm{Z}$. For $\bm{q}$, we concatenate the input point cloud with the diffusion sample $\bm{p}_{t}$ along the feature channel dimension, enabling the transformer architecture to learn explicit correspondences between the noisy canonical-frame partial cloud and the world-space partial cloud. For the geometric latent $\bm{Z}$, we apply a cross-attention mechanism to inject it into the point diffusion transformer. This approach ensures that the model effectively incorporates spatial and geometric relations. Additionally, due to symmetry and replicated geometrical shapes, multiple valid $\bm{p}$ may exist for a given $\bm{q}$ and $\bm{Z}$. Our diffusion model addresses this by sampling multiple noise realizations and aggregating the resulting transformations, to select the most confident and coherent representations.
4.3. Iterative Generation Procedure